OpenStudy (anonymous):

if a coin is tossed 3 times, what is the probability of setting 3 tails?

OpenStudy (anonymous):

there are 8 possible outcomes, and since 8 is not such a big number we can write them all out

OpenStudy (anonymous):

its in a fraction though

OpenStudy (anonymous):

(h, h, h) (h, h, t) (h t h) (t, h, h) (h, t, t) (t, h, t) (t. t. h) (t,t,t)

OpenStudy (anonymous):

no there are 6 possible outcomes , 2 each time you flip

OpenStudy (anonymous):

of those 8 only one has all tails

OpenStudy (anonymous):

answer is therefore \[\frac{1}{8}\]

OpenStudy (anonymous):

two possible outcome for each, you do it three times, counting principle gives \(2^3=8\) outcomes, not \(2\times 3=6\)

OpenStudy (anonymous):

Nvm satellite73 is completly right, read it wrong sorry!